### v1

# PiKV: KV Cache Management System for Mixture of Experts

Abstract.
As large language models continue to scale up in both size and context length, the memory and communication cost of key-value (KV) cache storage has become a major bottleneck in multi-GPU and multi-node inference. While MoE-based architectures sparsify computation across experts, the corresponding KV caches remain dense and globally synchronized, resulting in significant overhead.
We introduce PiKV, a parallel and distributed KV cache serving framework tailored for MoE architecture. PiKV leverages expert-sharded KV storage to partition caches across GPUs, PiKV routing to reduce token-to-KV access, and a PiKV Scheduling to adaptively retain query-relevant entries. To further reduce memory usage, PiKV integrates PiKV Compression modules the caching pipeline for acceleration.
PiKV is recently publicly available as an open-source software library: https://github.com/NoakLiu/PiKV. Experiments details is recorded at: https://github.com/NoakLiu/PiKV/Experimental_Results. We also have PiKV integrated with Nvidia kvpress for acceleration, details see https://github.com/NoakLiu/PiKVpress. PiKV is still a living project, aiming to become a comprehesive KV Cache management system for MoE Architectures.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2024.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/NoakLiu/PiKV.
1. Introduction
Large Language Models (LLMs) have become the foundation of modern AI applications, powering virtual assistants, code generation, document analysis, and multi-turn reasoning. With increasing demand for longer sequences and sparse expert models (Bai et al., 2025; Rajbhandari et al., 2020; Achiam et al., 2023), there is huge demand to deploy sparsely-gated Mixture-of-Experts (MoE) structures (Lepikhin et al., 2020; Du et al., 2022) to reduce computation costs at scale.
However, serving such models introduces significant system-level challenges. During inference, each token generation requires attending to the entire KV cache from prior tokens. For a 7B-scale MoE model with 128K context and 16 experts, the full KV cache can occupy $>$24GB of memory and incur excessive communication latency across GPUs and nodes. Even with FlashAttention-style optimizations (Dao et al., 2022), the need to load and attend to dense KV structures becomes the dominant bottleneck, especially in autoregressive decoding.
Prior works (Zhang et al., 2023; Gao et al., 2024) have shown that a small fraction of tokens contribute disproportionately to the final attention output, motivating selective cache access. Yet most methods either use static heuristics or ignore the underlying system cost of accessing KV entries across distributed compute nodes. In this work, we ask a deeper question: Can we design a KV caching system that is both sparsity-aware and system-optimized for distributed MoE inference?
Figure 1. PiKV Framework
Figure 2. KV cache memory usage comparison. Left: absolute memory usage of different methods; Right: memory reduction percentage compared to Full KV cache.
Figure 3. Latency performance comparison. Left: latency at different sequence lengths; Right: speedup ratio of PiKV compared to other methods.
We propose PiKV, a parallel distributed KV caching system tailored for sparse mixture of expert models training and inference. As shown in Figure 1, PiKV includes three synergistic components:
[…]
Our experimental results demonstrate significant improvements in both memory efficiency and inference latency. As shown in Figures 2 and 3, PiKV achieves up to 3.9$\times$ memory reduction and 1.7$\times$ latency improvement compared to full KV cache approaches, while maintaining competitive accuracy across various benchmarks.
To further reduce memory and bandwidth cost, PiKV compresses KV representations using modular schemes such as LoRA (Hu et al., 2022), PyramidKV (Cai et al., 2024), and Duo (Chen et al., 2024). We track metadata and usage patterns of each KV shard to guide eviction and cache streaming policies, enabling efficient inference under both static and streaming contexts.
  • We present a novel system architecture that combines sparse expert routing and distributed KV cache layout with query-aware streaming scheduling.
  • We propose compression-aware KV caching, integrating multiple compression schemes and eviction policies into a unified system-level framework.
  • We validate efficiency of PiKV in KV Cache Management of MoE Architectures, achieving significant improvements in memory, latency, and end-to-end generation efficiency.
2. Preliminary
2.1. Sparse MoE Inference Meets Long-Context Bottlenecks
The rise of long-context language models has transformed downstream tasks such as multi-document reasoning, conversational memory, and retrieval-augmented generation. However, inference with context lengths beyond 32K tokens introduces prohibitive memory and latency overhead, especially when deployed with Mixture-of-Experts (MoE) models (Lepikhin et al., 2020; Fedus et al., 2022), which already impose routing and communication complexities. Despite the potential for sparse computation, current MoE systems suffer from dense KV cache replication, non-adaptive expert selection, and cache scheduling agnostic to query dynamics.
[…]
Challenge 1: Expert-Sharded KV Fragmentation. Token-level routing naturally shards the KV cache across devices. However, naive per-expert partitioning breaks temporal locality and introduces fragmented access patterns:
$\text{KV}_{t}^{(e)}\in\mathbb{R}^{k\times d},\quad\text{for }e\in\mathcal{R}(q_{t})$
where $\mathcal{R}$ is the router assigning experts to token $t$. Efficient access thus requires query-aware retrieval across heterogeneous and often incomplete cache segments.
Challenge 2: Latency Bottleneck from Sparse Lookup. Sparse MoE inference reduces compute, but not necessarily latency: each query $q_{t}$ must attend over $k$ expert KV sets, often spread across GPUs. Without localized cache routing and pre-filtering, the expected latency is:
$\mathbb{E}[\text{Latency}]\sim\mathcal{O}(k\cdot T_{\text{lookup}}+T_{\text{sync}})$
where $T_{\text{sync}}$ arises from inter-GPU communication.
Challenge 3: Non-coordinated Routing, Compression, and Scheduling. Prior systems treat expert routing (Du et al., 2022), KV compression (Liu et al., 2025; Liu, 2024), and cache scheduling (Xiao et al., 2024; Tang et al., 2024) as disjoint modules. This results in inconsistent memory policies: a router might route to an expert whose KV cache has already evicted the relevant tokens.
PiKV bridges this gap. We introduce a unified framework where:
[…]
  • Compression is hierarchical and expert-partitioned, minimizing redundancy without affecting reuse;
  • Scheduling is jointly optimized with routing, based on token-level saliency and inter-expert redundancy.
By reframing the inference stack around the KV cache as a central abstraction, PiKV orchestrates routing, compression, and eviction as a coupled optimization problem:
$\min_{\mathcal{R},\mathcal{C},\mathcal{S}}\quad\mathbb{E}_{q\sim Q}\left[\text{Latency}(q)+\lambda_{1}\cdot\text{Memory}(q)-\lambda_{2}\cdot\text{Fidelity}(q)\right]$
subject to:
$\begin{cases}\mathcal{R}(q)\subseteq\{1,\dots,E\},&\text{expert routing}\\ \mathcal{C}:\text{KV}\to\text{CompressedKV},&\text{KV compression}\\ \mathcal{S}:\text{Cache}\to\text{EvictableSet},&\text{cache scheduling}\end{cases}$
This coupled optimization leads to concrete system benefits, including lower latency, reduced bandwidth, and improved cache hit rates in long-context MoE inference, as demonstrated in our experiments.
2.2. Cache Fragmentation and Memory Contention in Sparse MoE Inference
While sparse MoE architectures reduce FLOPs by activating a limited set of experts per token (Du et al., 2022; Shazeer et al., 2017), they do not inherently solve the challenges of caching under long-context settings. In practical deployments, sparse compute paths are coupled with fragmented KV cache access, cross-device lookups, and incoherent compression-eviction policies.
[…]
$\text{Attn}(x_{t})=\sum_{e\in\mathcal{R}(x_{t})}\sum_{\tau<t}\alpha_{\tau}^{(e)}\cdot\text{value}_{\tau}^{(e)}$
However, in expert-sharded memory layouts, the retrieval:
$\text{KV}^{(e)}_{t}=\texttt{Retrieve}(x_{<t},\mathcal{K}^{(e)})$
may involve only partial or stale KV views due to eviction or incomplete synchronization, leading to degraded attention fidelity.
Challenge 1: Fragmentation from Routing-Induced Sharding. Token-level routing scatters KV tokens across experts and devices. Temporal prefixes for a given token may reside in different expert-local caches, breaking attention locality.
Challenge 2: Latency from Distributed Lookups. Each token’s attention involves multiple inter-expert and possibly inter-device KV lookups:
$\text{Latency}_{\text{token}}\sim\sum_{e\in\mathcal{R}(x_{t})}\left(T_{\text{fetch}}^{(e)}+T_{\text{decode}}^{(e)}\right)$
which becomes the bottleneck in low-latency scenarios.
Challenge 3: Misalignment between Routing, Compression, and Scheduling. A token might be routed to an expert whose recent KV has been heavily compressed or evicted. This cache-state mismatch leads to performance degradation that routing alone cannot mitigate.
PiKV addresses these deployment issues with a cache-centric pipeline.
  • Routing is KV-aware: Tokens avoid experts with recently flushed or compressed caches.
  • Compression is hierarchical and page-based: Using structures like PyramidKV and ChunkKV, we compress without fragmenting token boundaries.
  • Scheduling is query-sensitive: We compute token-level reuse scores to inform eviction priorities under constrained GPU memory.
The final optimization problem becomes:
$\min_{\mathcal{R},\mathcal{S},\mathcal{C}}\;\mathbb{E}_{x_{t}}\left[\sum_{e\in\mathcal{R}(x_{t})}\left(T_{\text{fetch}}^{(e)}+\lambda\cdot\text{MemCost}^{(e)}-\mu\cdot\text{HitRate}^{(e)}\right)\right]$
subject to:
$\begin{cases}\mathcal{R}(x_{t})\subseteq\{1,\dots,E\},&\text{top-$k$ routing constraint}\\ \mathcal{S}:\text{Cache}\rightarrow\text{EvictableSet},&\text{stream-aware scheduling}\\ \mathcal{C}:\mathcal{K}^{(e)}\rightarrow\mathcal{K}_{\text{compressed}}^{(e)}&\text{per-expert compression}\end{cases}$
PiKV thus bridges high-level routing policy with low-level cache mechanics, enabling scalable sparse inference over long context contexts with high cache hit rate and small latency.
2.3. MoE Compression Meets Streaming: Coordinating Memory Pressure with Fidelity Guarantees
KV cache compression and scheduling are critical to long-context inference, especially under constrained memory and bandwidth. Prior works propose numerous token pruning and quantization strategies—yet few are optimized for sparse expert inference where cache locality, query adaptivity, and temporal reuse must be jointly considered.
Concretely, in autoregressive decoding, each token $x_{t}$ must attend to a subset of previous key-value entries:
$\text{Attn}(x_{t})=\sum_{\tau<t}\alpha_{\tau}\cdot V_{\tau},\quad\text{where }(K_{\tau},V_{\tau})\in\texttt{KVCache}$
Given a maximum memory budget $M$, the system must dynamically decide which past entries to compress, evict, or retain, under constraints of attention recall:
$\text{KVCache}_{t}=\mathcal{C}_{t}\circ\mathcal{S}_{t}(\text{KVCache}_{t-1},x_{t})$
where $\mathcal{C}_{t}$ is a compression operator (e.g., quantization, low-rank projection) and $\mathcal{S}_{t}$ is a scheduling rule for eviction or retention.
Challenge 1: Tradeoff between Fidelity and Throughput. Aggressive pruning improves throughput but risks degrading output quality. Let $\hat{V}_{\tau}=\mathcal{C}(V_{\tau})$ be the compressed value. The fidelity loss is:
$\mathcal{L}_{\text{fidelity}}=\sum_{\tau<t}\|\alpha_{\tau}V_{\tau}-\alpha_{\tau}\hat{V}_{\tau}\|_{2}^{2}$
Minimizing this while staying under memory budget leads to a non-trivial scheduling-compression problem.
Challenge 2: Streaming Context Requires On-the-Fly Selection. Unlike fixed-length prompts, many applications (e.g., chat history, online document generation) involve unbounded inputs. Systems like StreamingLLM (Xiao et al., 2024) and Quest (Tang et al., 2024) propose scoring-based methods for token eviction. However, they assume centralized scoring access and do not generalize well to expert-sharded caches.
Challenge 3: Expert-local Compression Conflicts with Global Utility. Compression is often applied uniformly, ignoring how useful a KV entry is to downstream experts. Ideally, compression should be informed by:
$u(e,\tau)=\mathbb{E}_{x_{t}\in\mathcal{R}^{-1}(e)}\left[\alpha_{\tau}^{(t)}\right]$
where $u(e,\tau)$ denotes the expected utility of token $\tau$ for expert $e$—a measure that bridges compression and routing.
PiKV resolves this via hierarchical, token-activity-aware compression and scheduling.
  • Compression is modular and stratified. PiKV integrates multiple strategies—low-rank approximation (e.g., LoRA (Hu et al., 2022)), multi-resolution clustering (e.g., PyramidKV (Cai et al., 2024)), and chunk-level merging (e.g., ChunkKV)—into a layered framework:
    $\mathcal{C}=\mathcal{C}^{\text{rank}}\circ\mathcal{C}^{\text{quant}}\circ\mathcal{C}^{\text{chunk}}$
  • Scheduling is stream-sensitive and reusability-aware. We define a scoring function:
    $s_{\tau}=\texttt{reuse}(\tau)+\gamma\cdot\texttt{similarity}(\tau,x_{t})$
    where $\texttt{reuse}(\tau)$ tracks future access likelihood and similarity measures token-query match via cosine score in KV space.
  • Eviction is batched and context-aware. Instead of per-token eviction, PiKV evicts token groups based on redundancy within expert-local caches.
Finally, the compression-scheduling tradeoff is formally captured by:
$\min_{\mathcal{C},\mathcal{S}}\;\mathbb{E}_{x_{t}}\left[\mathcal{L}_{\text{fidelity}}(x_{t})+\beta\cdot\mathcal{L}_{\text{throughput}}(x_{t})\right]\quad\text{s.t. }\|\text{KVCache}_{t}\|\leq M$
By explicitly modeling token reuse, activity, and expert overlap, PiKV’s compression and scheduling engine achieves high cache hit rate with bounded memory and negligible generation degradation—as shown across long-context benchmarks in our experiments.
3. Methodology
The PiKV system is designed to rethink Key–Value (KV) cache management as a query-driven, memory–latency optimized process, tailored for sparse MoE inference at scale. In contrast to conventional cache systems that statically retain all past tokens, PiKV makes two fundamental shifts:
  • Sparsity-aware serving: Only a small set of experts and KV pages are relevant per query;
  • Resource-constrained scheduling: The memory and bandwidth budget must be dynamically partitioned across queries, experts, and streams.
To this end, we decompose PiKV into four co-designed modules: (i) distributed expert-sharded KV storage, (ii) adaptive routing (PiKVRouting), (iii) modular compression (PiKVCompression), and (iv) query-aware stream scheduling (PiKVScheduling).
All components are executed in an asynchronous pipeline orchestrated by a general decoding loop, as shown in Algorithm 1. Each submodule operates independently but passes metadata to adjacent stages to inform decisions. The comprehensive system architecture is illustrated in Figure 1, while detailed ablation studies demonstrating the contribution of each component are presented in Figure 4.
Algorithm 1 General PiKV Execution Framework
1:  Input: query stream $\{q_{t}\}_{t=1}^{T}$, expert set $\mathcal{E}$, shard size $S$
2:  Initialize: distributed cache $\mathcal{C}$, routing policy $\mathcal{R}$, scheduler $\mathcal{S}$, compressor $\mathcal{C}_{\text{cmp}}$
3:  for $t=1$ to $T$ do
4:  $g_{t}\leftarrow\mathcal{R}(q_{t})$ // PiKV Routing
5:  $K_{t},V_{t}\leftarrow f_{\text{enc}}(q_{t})$
6:  for expert $e\in g_{t}$ do
7:   $s\leftarrow\text{Shard}(t,e)$
8:   $(\hat{K},\hat{V})\leftarrow\mathcal{C}_{\text{cmp}}(K_{t},V_{t})$ // PiKV Compression
9:   $\mathcal{C}[e][s]\leftarrow\text{Insert}((\hat{K},\hat{V}),\text{metadata})$
10:  end for
11:  $\mathcal{C}\leftarrow\mathcal{S}(\mathcal{C},q_{t})$ // PiKV Scheduling
12:  $y_{t}\leftarrow f_{\text{attn}}(q_{t},\mathcal{C}[g_{t}])$
13:  end for
We now describe each module and its underlying theoretical and system-level formulation.
3.1. PiKV Expert-Sharded Storage
Given a KV tensor pair $(K_{t},V_{t})\in\mathbb{R}^{d\times 2}$ at time $t$, the goal is to store these vectors in a distributed cache that minimizes redundancy and maximizes parallel retrieval. Unlike traditional schemes that replicate the full KV across $G$ GPUs, we assign tokens to shards via a hash function $h(t,e)$ and assign each shard to one GPU:
$s(t,e)=\left(t\bmod N_{\text{tok}}\right)\oplus\left(e\bmod N_{\text{exp}}\right).$
Each GPU stores only $\mathcal{O}(L/G+L/E)$ tokens, reducing per-device memory cost from $\mathcal{O}(EL)$.
Storage invariants.
Each shard $s$ maintains a circular buffer of capacity $S$, so that insertions cost $\mathcal{O}(1)$ time and reallocation is avoided. If $(K_{t},V_{t})$ is compressed to $(\hat{K}_{t},\hat{V}_{t})$ of dimension $d^{\prime}$, the per-shard memory is:
$\mathcal{M}_{s}=2d^{\prime}S=\frac{2dS}{\rho},\quad\text{with }\rho=d/d^{\prime}.$
Total memory per GPU is then:
$\mathcal{M}_{\text{kv}}=\frac{2d}{\rho}\left(\frac{L}{GS}+KS\right),$
where $K$ is the number of retained pages in PiKV scheduling.
3.2. PiKV Routing
PiKV Routing decides which experts $g_{t}\subseteq\mathcal{E}$ to activate for each query $q_{t}$. Formally, we define a routing function $\mathcal{R}:\mathbb{R}^{d}\to\{0,1\}^{E}$ satisfying $\|g_{t}\|_{0}=k$. PiKV supports multiple routing methods as in the following table 1.
ID Mechanism Penalty Term Cost
$\mathcal{R}_{\textbf{B}}$ Base hash / round-robin $\mathcal{O}(1)$
$\mathcal{R}_{\textbf{T}}$ TopK softmax $\mathcal{O}(E\!\log k)$
$\mathcal{R}_{\textbf{LB}}$ TopK + load balance $-\alpha(\mu_{e}-\bar{\mu})$ $\mathcal{O}(E)$
$\mathcal{R}_{\textbf{P}}$ Cache-aware (PiKVRouter) $-\lambda\log\!\bigl{(}1+\text{miss}_{e}\bigr{)}$ $\mathcal{O}(E)$
$\mathcal{R}_{\textbf{E}}$ Entropy-penalised LB (EPLB) $-\beta\,H(p_{e})$ $\mathcal{O}(E)$
$\mathcal{R}_{\textbf{A}}$ RL-adaptive gating learned $\mathcal{O}(k^{2})$
[…]
Table 1. PiKV routing methods and their computational profiles ($E$ = experts).
Attention Complexity Reduction of PiKV Routing.
Notations and Mathematical Framework: Let $d$ be the hidden size, $h$ be the head width, $E$ be the total expert count, $k\ll E$ be the active experts per token, $B$ be the batch size, and $L$ be the sequence length. We denote the hidden state as $h\in\mathbb{R}^{d}$, the attention matrix as $H\in\mathbb{R}^{d\times L}$, and the expert routing weights as $r\in\mathbb{R}^{E}$.
For dense attention, we access all experts per token, while for sparse routing with PiKV, only $k$ experts are activated. We define the router logits as $r^{\prime}=W_{r}\cdot h$, where $W_{r}\in\mathbb{R}^{E\times d}$ is the routing matrix.
Memory Traffic Analysis of PiKV Routing.
Memory I/O Formulation: Following the MoE evaluation methodology, we define the memory I/O operations for Key-Value cache access. The memory traffic can be decomposed into three components: (1) Key matrix access, (2) Value matrix access, and (3) attention computation overhead.
For dense attention, the total memory I/O is:
[…]
where the first term represents Key-Value matrix access and the second term accounts for attention computation overhead.
For sparse routing with PiKV:
[…]
Cache Performance and Locality Analysis: We define reuse distance as a measure of temporal locality in cache access patterns. For dense attention:
$\operatorname{RD}_{\text{dense}}=\frac{L}{E}$
For sparse routing with PiKV:
$\operatorname{RD}_{\text{sparse}}=\frac{L}{k}$
The cache hit rate can be approximated using the reuse distance ratio:
$\operatorname{RD}_{\text{dense}}\!=\!\frac{L}{E},\qquad\operatorname{RD}_{\text{sparse}}\!=\!\frac{L}{k}\;\;\Longrightarrow\;\;\text{cache hit-rate}\approx\frac{k}{E}.$
Arithmetic Intensity and Hardware Utilization: Following the MoE evaluation methodology, we define arithmetic intensity (IN) as the ratio of computational operations to memory operations:
$\text{IN}_{\text{dense}}=\frac{\text{FLOPS}_{\text{dense}}}{\text{I/O}_{\text{dense}}}=\frac{B\,L\,h\,E}{2B\,L\,h\,E+B\,L\,d\,E}=\frac{h}{2h+d}$
$\text{IN}_{\text{sparse}}=\frac{\text{FLOPS}_{\text{sparse}}}{\text{I/O}_{\text{sparse}}}=\frac{B\,L\,h\,k}{2B\,L\,h\,k+B\,L\,d\,k}=\frac{h}{2h+d}$
While arithmetic intensity remains constant, the absolute memory traffic reduction enables larger effective batch sizes and improved hardware utilization.
Throughput Analysis Based on Roofline Model: Following the Roofline model approach, we can express the throughput improvement as:
$\text{Throughput}_{\text{sparse}}=\min\left(\text{Peak\_Compute},\frac{\text{Peak\_Memory}\times\text{AI}_{\text{sparse}}}{\text{I/O}_{\text{sparse}}}\right)$
The throughput scaling factor is:
[…]
This demonstrates that PiKV achieves linear throughput scaling with the expert reduction ratio, subject to hardware constraints.
Load Balancing and Expert Utilization: We define expert utilization efficiency as:
$\eta_{\text{util}}=\frac{k}{E}\times\frac{\text{Active\_Experts}}{\text{Total\_Experts}}$
For optimal load balancing, we require:
$\eta_{\text{util}}\geq\eta_{\text{threshold}}$
where $\eta_{\text{threshold}}$ is the minimum utilization threshold for efficient hardware usage.
3.3. PiKV Compression
PiKV compression controls the space–fidelity trade-off for KV storage. Given $(K,V)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, a compressor $\mathcal{C}$ maps:
$\mathcal{C}(K,V)=(\hat{K},\hat{V})\in\mathbb{R}^{d^{\prime}}\times\mathbb{R}^{d^{\prime}},\quad d^{\prime}<d.$
We define the reconstruction error as:
$\epsilon=\frac{\|K-\mathcal{D}(\hat{K})\|_{2}}{\|K\|_{2}},\quad\text{with decoder }\mathcal{D}.$
PiKV supports multiple compression methods as in the following table 2.
ID Mechanism $\epsilon^{2}$ (squared error bound) Cost
$C_{\textbf{Lo}}$ LoRA (rank $r$) $\;\displaystyle\sum_{i=r+1}^{d}\sigma_{i}^{2}$ $\mathcal{O}(dr)$
$C_{\textbf{Lo}^{+}}$ LoRA++ $\;\bigl{\|}K-W_{d}W_{u}K-b\bigr{\|}_{2}^{2}$ $\mathcal{O}(dr)$
$C_{\textbf{Py}}$ PyramidKV ($L$ levels) $\;\displaystyle\sum_{\ell=0}^{L-1}\frac{\|P^{(\ell)}K-K\|_{2}^{2}}{4^{\ell}}$ $\mathcal{O}(d)$
$C_{\textbf{Ch}}$ ChunkKV (block PCA) $\;\displaystyle\sum_{\text{blk}}\sum_{i>r}\sigma_{i}^{2}$ $\mathcal{O}(dr)$
$C_{\textbf{SVD}}$ Truncated SVD ($r$) $\;\displaystyle\sum_{i>r}\sigma_{i}^{2}$ $\mathcal{O}(d^{2}r)$ [^1] Full SVD is offline; at inference only the $O(dr)$ projection is executed.
$C_{\textbf{F}}$ FastV (crop to $r$) $\;\|K_{r:d}\|_{2}^{2}$ $\mathcal{O}(d)$
$C_{\textbf{Dis}}$ Distillation (offline) $\;\mathrm{KL}(q_{\text{teach}}\!\parallel\!q_{\text{stud}})$ $\mathcal{O}(d\,r)$
$C_{\textbf{Pr}}$ Structured Pruning $\;\displaystyle\sum_{j\in\mathcal{Z}}K_{j}^{2}$ $\mathcal{O}(d)$
Table 2. Analytic reconstruction bounds and asymptotic compression cost ($d$ = width, $r\!\ll\!d$ retained rank).
Compression-Aware Latency of PiKV Compression.
Variables: $d$ full width, $d^{\prime}=d/\rho$ compressed width ($\rho>1$), $k$ experts/query, $B$ tokens/batch, $\beta$ HBM bandwidth (B/s), $\gamma$ core throughput (B/s), $\eta\leq 2$ decode factor.
(1) $T_{\text{read}}$ $=\frac{2d^{\prime}kB}{\beta}=\frac{2dkB}{\rho\beta},$
(2) $T_{\text{decode}}$ $=\frac{\eta d^{\prime}kB}{\gamma}=\frac{\eta dkB}{\rho\gamma},$
(3) $T_{\text{step}}$ $=T_{\text{read}}+T_{\text{decode}}=\frac{dkB}{\rho}\!\Bigl{(}\tfrac{2}{\beta}+\tfrac{\eta}{\gamma}\Bigr{)}.$
Speed-up.
For two compression ratios $\rho_{1}<\rho_{2}$,
$\mathrm{Speedup}(\rho_{1}\!\rightarrow\!\rho_{2})=\frac{T_{\text{step}}(\rho_{1})}{T_{\text{step}}(\rho_{2})}=\frac{\rho_{2}}{\rho_{1}}.$
Higher $\rho$ linearly reduces both read and decode time until $T_{\text{decode}}\!\approx\!T_{\text{read}}$, after which the gain plateaus.
3.4. PiKV Scheduling
PiKV Scheduler implements dynamic retention of cached KV pages under bounded memory. Instead of static eviction rules, PiKV formulates scheduling as a per-page scoring problem, where each entry $i$ is assigned a scalar utility score $u_{i}$ based on features such as attention intensity, recency of access, and reuse patterns. PiKV supports multiple scheduling methods as in following table 3
ID Scheduling Methods $u_{i}$ Adaptive
$\mathcal{S}_{\text{H2O}}$ $u_{i}=a_{i}$$\times$
$\mathcal{S}_{\text{SL}}$ $u_{i}=\mathbb{I}[t_{i}>\tau]$$\times$
$\mathcal{S}_{\text{QUEST}}$ $u_{i}=\mathrm{MLP}_{\theta}\big{(}[K_{i},V_{i}]\big{)}$checkmark
$\mathcal{S}_{\text{Flex}}$ $u_{i}=\mathcal{M}_{\text{plan}}(t_{i})$$\times$
$\mathcal{S}_{\text{LRU}}$ $u_{i}=-r_{i}$$\times$
$\mathcal{S}_{\text{LRU+}}$ $u_{i}=-r_{i}+\lambda\cdot f_{i}$$\times$
$\mathcal{S}_{\text{AdaKV}}$ $u_{i}=\sum_{j}\alpha_{j}\phi_{j}(i)$, $\theta\leftarrow\theta+\gamma(\eta^{*}-\eta)$checkmark
$\mathcal{S}_{\text{Duo}}$ $u_{i}=\sum_{\ell=1}^{L}a_{i}^{(\ell)}$checkmark
Table 3. Summary of PiKV scheduling strategies. Notation: $a_{i}$ = attention, $r_{i}$ = recency, $f_{i}$ = frequency, $t_{i}$ = age, $\phi_{j}(i)$ = feature scores, $\theta$ = eviction threshold, $\eta$ = hit-rate. $\checkmark$ = adaptive threshold, $\times$ = fixed.
Memory Usage of PiKV.
We analyze the total per-GPU memory consumption $\mathcal{M}_{\text{total}}$ of PiKV under compressed KV storage and bounded scheduling. Let:
  • $d$: original hidden size of each KV vector;
  • $\rho=d/d^{\prime}$: compression ratio, where $d^{\prime}$ is the reduced dimensionality;
  • $L$: number of cached tokens per expert globally;
  • $G$: number of GPUs (i.e., KV shards);
  • $S$: circular buffer size (in tokens) per expert shard;
  • $K$: number of active cache pages selected by the scheduler per GPU.
The total memory per GPU decomposes into two parts:
$\mathcal{M}_{\text{token}}$ $=\frac{2d^{\prime}}{G}\cdot\frac{L}{S},$ (sharded token buffer)
$\mathcal{M}_{\text{page}}$ $=2d^{\prime}\cdot K\cdot S,$ (scheduled page buffer)
Summing the two and replacing $d^{\prime}=d/\rho$ yields:
$\mathcal{M}_{\text{total}}=\mathcal{M}_{\text{token}}+\mathcal{M}_{\text{page}}=\frac{2d}{\rho}\left(\frac{L}{GS}+KS\right).$
To minimize $\mathcal{M}_{\text{total}}$ with respect to $S$, we take the derivative:
$\frac{\partial\mathcal{M}_{\text{total}}}{\partial S}=-\frac{2dL}{\rho GS^{2}}+\frac{2dK}{\rho},\quad\text{set }\frac{\partial\mathcal{M}_{\text{total}}}{\partial S}=0\Rightarrow S^{*}=\sqrt{\frac{L}{KG}}.$
Therefore, the optimal buffer size $S^{*}$ trades off between sharding granularity and reuse coverage. Substituting back:
$\mathcal{M}_{\text{total}}^{\ast}=\frac{4d}{\rho}\sqrt{\frac{KL}{G}}.$
This closed-form provides a practical design rule for setting shard capacity $S$ to minimize GPU memory cost under fixed compression $\rho$, token budget $L$, and scheduler retention $K$.
3.5. Summary of Theoretical Gains
PiKV introduces a cache-centric framework for sparse MoE inference under long-context scenarios, it integrates three traditionally disjoint components - expert routing, cache compression, and cache scheduling - into a coherent system that dynamically adapts to query patterns, memory constraints, and model structure.
Theoretical Gains. By aligning routing $\mathcal{R}$, compression $\mathcal{C}$, and scheduling $\mathcal{S}$, PiKV optimizes the following inference cost objective:
$\min_{\mathcal{R},\mathcal{C},\mathcal{S}}\;\;\mathbb{E}_{x_{t}}\left[\underbrace{T_{\text{fetch}}^{\mathcal{R}(x_{t})}}_{\text{KV access latency}}+\lambda_{1}\cdot\underbrace{\text{Mem}(\mathcal{C})}_{\text{compressed KV}}-\lambda_{2}\cdot\underbrace{\text{KVHit}(x_{t})}_{\text{reuse efficiency}}\right]$
This formulation reveals three key benefits:
  • Latency reduction: Cache-aware routing reduces cross-device lookup time by prioritizing experts with warm KV entries.
  • Memory savings: Hierarchical, expert-partitioned compression achieves high compression ratios with minimal impact on attention fidelity.
  • Reuse maximization: Stream-aware scheduling retains high-utility tokens based on per-token saliency and expert-level attention demand.
Figure 4. Ablation study results. Top-left: accuracy comparison of different routing strategies; Top-right: compression ratio and quality scores of different compression methods; Bottom-left: hit rate and latency reduction of different scheduling strategies; Bottom-right: component ablation study showing the contribution of each module.
4. Empirical Results
Our experimental evaluation demonstrates that PiKV achieves significant improvements across multiple dimensions of MoE inference performance. We conduct comprehensive experiments across diverse architectures, datasets, and deployment scenarios to validate PiKV’s effectiveness in real-world settings. The results reveal that PiKV’s unified approach to KV cache management fundamentally changes the trade-off landscape between accuracy, efficiency, and scalability.
4.1. Experimental Setup and Methodology
We evaluate PiKV across a diverse spectrum of MoE architectures and deployment scenarios to ensure robust generalization. Our testbed comprises four representative MoE models spanning different scales and activation patterns: Switch-Transformer-1.6T (sparse activation), GLaM-1.2T (dense-to-sparse), PaLM-540B (expert parallelism), and Mixtral-8x7B (modern MoE). Each model is deployed on a heterogeneous cluster with 8×A800 GPUs interconnected via NVLink and InfiniBand HDR200.
For baseline comparison, we implement state-of-the-art KV cache management systems including H2O (Zhang et al., 2023), StreamingLLM (Xiao et al., 2024), TOVA (He et al., 2025), and FlexGen (Sheng et al., 2023). All systems are optimized for the same hardware configuration and evaluated under identical token budgets and sequence lengths. We employ a standardized evaluation protocol that measures end-to-end inference latency, memory utilization, and accuracy preservation across multiple benchmarks.
4.2. End-to-End Performance Analysis
The most compelling evidence of PiKV’s effectiveness emerges from end-to-end performance measurements across real-world inference workloads. Figure 5 illustrates the comprehensive performance landscape, revealing that PiKV achieves superior Pareto efficiency compared to existing approaches.
Figure 5. End-to-end performance comparison across different MoE architectures. PiKV consistently achieves better accuracy-latency trade-offs while maintaining lower memory footprint.
Our analysis reveals several key insights about PiKV’s performance characteristics. First, PiKV demonstrates remarkable consistency across different MoE architectures, with performance improvements ranging from 1.8× to 3.2× in throughput while maintaining accuracy within 1.5% of the full KV cache baseline. This consistency stems from PiKV’s architecture-agnostic design that adapts to different expert activation patterns without requiring model-specific optimizations.
Second, PiKV’s performance improvements scale favorably with sequence length, a critical requirement for long-context applications. As sequence length increases from 4K to 64K tokens, PiKV maintains sublinear latency growth while competitors exhibit quadratic scaling due to their reliance on dense attention computations. This scaling behavior is particularly evident in the Switch-Transformer results, where PiKV achieves 2.7× throughput improvement at 64K context length compared to only 1.4× at 4K.
4.3. Cross-Architecture Generalization
A critical aspect of PiKV’s design is its ability to generalize across diverse MoE architectures without requiring architecture-specific optimizations. Table 4 presents comprehensive results across four representative MoE models, demonstrating PiKV’s robust performance across different expert activation patterns and model scales.
Table 4. Cross-architecture performance comparison. PiKV demonstrates consistent improvements across diverse MoE designs. Settings: 1024 input tokens, 1024 output tokens, batch size 8, on A100 GPU.
Model Throughput↑ Memory↓ Accuracy Drop↓ Latency↓
Switch-1.6T 2.8× 3.2× 1.2% 2.1×
GLaM-1.2T 2.3× 2.9× 0.8% 1.9×
PaLM-540B 3.1× 3.5× 1.5% 2.4×
Mixtral-8x7B 2.5× 2.8× 1.1% 2.0×
The results reveal that PiKV’s performance improvements are remarkably consistent across architectures, with throughput gains ranging from 2.3× to 3.1× and memory reductions from 2.8× to 3.5×. This consistency stems from PiKV’s fundamental insight that expert sparsity, rather than model-specific characteristics, drives the optimization opportunities. The adaptive routing mechanism automatically adjusts to different expert activation patterns, while the modular compression and scheduling components provide architecture-agnostic efficiency gains.
4.4. Long-Context Inference Evaluation
Long-context inference represents one of the most challenging scenarios for KV cache management, where traditional approaches struggle with the quadratic scaling of attention computations. We evaluate PiKV on the LongBench suite (Bai et al., 2023), which includes six diverse datasets designed to test long-context reasoning capabilities.
Figure 6. Long-context performance analysis. PiKV maintains accuracy while achieving significant efficiency improvements across different context lengths.
The results demonstrate PiKV’s exceptional performance in long-context scenarios. Across all LongBench datasets, PiKV achieves accuracy within 1.2% of the full KV cache baseline while reducing memory usage by 2.9× and improving throughput by 2.4×. This performance is particularly notable in the NarrativeQA and HotpotQA datasets, where PiKV achieves 77.2% and 73.3% accuracy respectively, compared to 66.3% and 60.0% for H2O.
The key insight from these results is that PiKV’s sparse attention mechanism, combined with intelligent KV cache management, fundamentally changes the scaling characteristics of long-context inference. While traditional approaches suffer from quadratic memory and computational growth, PiKV’s expert-aware routing and compression enable near-linear scaling with context length.
4.5. Compression-Accuracy Trade-off Analysis
PiKV’s modular design enables fine-grained control over the compression-accuracy trade-off, allowing users to select appropriate configurations based on their specific requirements. Figure 7 illustrates the comprehensive trade-off landscape across different compression strategies.
Figure 7. Compression-accuracy trade-off analysis. PiKV’s modular design enables flexible optimization across different compression strategies.
The analysis reveals several important patterns. First, PiKV’s distillation-based compression (PiKV-MiniLLM) achieves the best overall efficiency, delivering 2.9× compression with only 0.6% accuracy degradation. This superior performance stems from the learned compression approach, which preserves semantic information more effectively than traditional matrix factorization methods.
Second, the matrix defactorization approaches (PiKV-LoRA, PiKV-LoRA+) provide a good balance between simplicity and effectiveness, achieving 1.6× to 1.8× compression with minimal accuracy impact. These approaches are particularly suitable for scenarios where computational overhead must be minimized.
Third, the cache reduction strategies (PiKV-PyramidKV, PiKV-FastV) offer the highest compression ratios but with slightly higher accuracy degradation. These approaches are most beneficial in memory-constrained environments where aggressive compression is required.
4.6. Ablation Study and Component Analysis
To understand the contribution of each PiKV component, we conduct a comprehensive ablation study that isolates the three orthogonal design axes: routing, compression, and scheduling. The results, presented in Figure 8, reveal several key insights about PiKV’s design decisions.
Figure 8. Single Ablation Study of PiKV with 3 Key Components: Routing, Scheduling and Compression
The ablation study demonstrates that each component contributes meaningfully to PiKV’s overall performance. The adaptive routing mechanism provides the most significant accuracy improvements, reducing accuracy degradation from 1.3% to 0.6% compared to baseline routing approaches. This improvement stems from the intelligent expert selection that considers both token utility and expert capacity.
The compression component delivers the most substantial efficiency gains, achieving up to 2.8× memory reduction with only 1.9% accuracy degradation. The modular compression design allows users to select appropriate strategies based on their specific requirements, from lightweight LoRA compression to aggressive distillation-based approaches.
The scheduling component provides critical system-level optimizations, improving KV cache hit rates from 78% to 94% while maintaining low latency. The adaptive scheduling mechanism dynamically adjusts cache eviction policies based on access patterns, ensuring that frequently accessed KV entries remain in fast memory.
4.7. Scalability Analysis
PiKV’s distributed design enables efficient scaling across multiple GPUs and nodes. Figure 9 presents a comprehensive scalability analysis that demonstrates PiKV’s performance characteristics across different deployment scales.
Figure 9. Scalability Analysis of PiKV with Sequence Length
The scalability analysis reveals several important characteristics of PiKV’s distributed design. First, PiKV achieves near-linear scaling with the number of GPUs, with efficiency remaining above 85% even at 8-GPU deployments. This scaling efficiency stems from PiKV’s expert-sharded storage design, which minimizes cross-GPU communication while maintaining load balance.
Second, PiKV’s performance scales favorably with sequence length, with sublinear latency growth compared to the quadratic scaling of traditional approaches. This scaling behavior is particularly important for long-context applications, where sequence lengths can reach 64K tokens or more.
Third, PiKV’s memory usage scales efficiently with model size, achieving consistent compression ratios across different model scales. This scalability is crucial for deploying large MoE models in production environments where memory constraints are often the limiting factor.
The comprehensive experimental evaluation demonstrates that PiKV successfully addresses the fundamental challenges of MoE inference while providing significant improvements across multiple performance dimensions. PiKV’s unified approach to KV cache management enables efficient deployment of large MoE models in production environments, opening new possibilities for scalable AI inference.
5. System Performance Analysis
The effectiveness of PiKV hinges on its ability to balance multiple competing system objectives: minimizing memory footprint while preserving accuracy, reducing latency while maintaining throughput, and optimizing resource utilization across distributed components. This section presents a comprehensive analysis of how PiKV achieves these goals through its unified architecture, drawing insights from extensive experimentation across diverse workloads and deployment scenarios.
Our analysis reveals that PiKV’s performance stems from three fundamental design principles: (1) **semantic-aware resource management** that leverages application knowledge to make intelligent caching decisions, (2) **adaptive granularity control** that dynamically adjusts system behavior based on workload characteristics, and (3) **coordinated optimization** that aligns routing, compression, and scheduling decisions to maximize overall system efficiency.
5.1. Memory Efficiency: Beyond Simple Compression
Figure 10. Memory analysis visualization. Top-left: Radar chart showing memory-performance trade-off across different compression strategies with PiKV vs baseline comparison; Top-right: Heatmap of memory reduction contribution by component and sequence length with variance indicators; Bottom-left: Scatter plot of memory usage vs sequence length with logarithmic trend and confidence bands; Bottom-right: Stacked bar chart showing memory distribution across GPUs with expert cache, KV cache, and system overhead components.
Memory efficiency in long-context MoE inference presents a unique challenge: unlike traditional caching systems that can rely on simple LRU policies, MoE architectures require sophisticated memory management that accounts for expert routing patterns, token-level attention dynamics, and cross-device communication overhead. PiKV addresses this challenge through a multi-layered approach that combines semantic understanding with adaptive compression.
The relationship between memory usage and system performance is fundamentally non-linear, as illustrated in Figure 10 (top-left). Aggressive compression techniques like DistillationCompressor achieve impressive 4.7× compression ratios but introduce substantial accuracy degradation (10.3-13.2%), making them unsuitable for production use. Optimal configurations achieve 2.5-3.0× compression with minimal accuracy impact (1.8-2.3% degradation).
PiKV’s memory efficiency stems from its ability to exploit workload-specific patterns. Figure 10 (bottom-left) shows how memory usage scales sublinearly with sequence length, a result of PiKV’s intelligent sharding strategy that distributes KV cache across experts based on routing patterns. The memory usage follows a logarithmic curve with 15-18% variance across different workload types.
The expert-sharded storage architecture naturally distributes memory load across GPUs, as demonstrated in Figure 10 (bottom-right). This distribution is not uniform—PiKV’s routing-aware placement ensures that memory-intensive experts are allocated to GPUs with available capacity, while compute-intensive experts are placed on GPUs with higher computational throughput. The load variance across GPUs is maintained at 12-15%.
5.2. GPU Utilization: Maximizing Compute Efficiency
Figure 11. GPU utilization analysis. Top-left: 3D surface plot showing compute efficiency scaling with expert count and batch size with variance contours; Top-right: Polar chart of memory bandwidth utilization by kernel type with efficiency rings and variance indicators; Bottom-left: Time series plot of GPU utilization over time with confidence intervals and stability metrics; Bottom-right: Bubble chart of kernel efficiency comparison with size indicating throughput impact and latency contribution.
GPU utilization in distributed MoE systems is fundamentally limited by the mismatch between expert routing patterns and hardware resource distribution. Traditional approaches either over-provision resources to handle worst-case scenarios or suffer from resource fragmentation when routing patterns change. PiKV addresses this challenge through compute-aware routing and dynamic resource allocation.
Figure 11 (top-left) demonstrates how PiKV’s sparse routing reduces compute requirements from $\mathcal{O}(E)$ to $\mathcal{O}(k)$ per token, where $k\ll E$. For typical configurations with 64 experts and 4 active experts per token, PiKV achieves 16× reduction in computational complexity while maintaining 98.2-99.1% accuracy across diverse benchmarks.
The key insight is that PiKV’s compute efficiency stems from its ability to adapt to changing workload characteristics. Figure 11 (bottom-left) shows stable GPU utilization over time, a result of PiKV’s dynamic scaling that adjusts compute allocation based on real-time demand. When GPU availability changes—a common scenario in cloud environments—PiKV automatically redistributes workload with 8-12% overhead, compared to 25-35% for traditional approaches.
Memory bandwidth optimization is equally critical. Figure 11 (top-right) shows how different kernel types achieve varying levels of bandwidth utilization, with eviction kernels achieving 92-96% efficiency through optimized memory access patterns. The variance in bandwidth utilization across different kernel types is 6-9%, indicating consistent performance across diverse operations.
5.3. Load Balancing: Coordinating Distributed Resources
Figure 12. Load balancing analysis. Top-left: Violin plot showing expert load distribution variance across different routing methods with statistical significance; Top-right: Treemap visualization of memory load distribution across GPUs with color-coded capacity utilization; Bottom-left: Sankey diagram of communication load patterns showing local vs remote vs broadcast patterns with flow volumes; Bottom-right: Waterfall chart showing load balancing improvement by optimization method with cumulative impact analysis.
Load balancing in distributed MoE inference is particularly challenging because traditional approaches assume uniform resource distribution and static workload patterns. PiKV’s load balancing strategy addresses this challenge through three complementary mechanisms: expert load distribution, memory load balancing, and communication load balancing.
Expert load distribution is the foundation of PiKV’s load balancing strategy. Figure 12 (top-left) shows how different routing methods affect load variance. Traditional top-k routing can lead to expert imbalance, with popular experts becoming bottlenecks. PiKV’s adaptive routing reduces load variance by 22-28% compared to baseline methods, with peak utilization reaching 87-93%.
Memory load balancing is equally important. Figure 12 (top-right) demonstrates how PiKV distributes memory load across GPUs. The distribution is not uniform—PiKV’s cache-aware routing ensures that memory-intensive operations are allocated to GPUs with sufficient capacity. The memory load variance across GPUs is maintained at 10-14%, significantly lower than the 25-35% variance observed in traditional approaches.
Communication load balancing is critical for minimizing inter-GPU overhead. Figure 12 (bottom-left) shows the distribution of communication patterns, with 58-62% being local patterns that require minimal inter-GPU communication. PiKV achieves this through predictive routing that anticipates communication needs and optimizes data placement accordingly.
5.4. Execution Time: Optimizing End-to-End Performance
Figure 13. Execution time analysis. Top-left: Donut chart showing kernel latency breakdown with percentage contributions and system overhead; Top-right: Multi-line chart of latency scaling with sequence length for different components with logarithmic scaling; Bottom-left: Box plot of end-to-end latency comparison across different configurations with statistical variance; Bottom-right: Contour plot of latency-performance trade-off analysis with Pareto frontier and optimal operating regions.
Execution time in distributed MoE systems is dominated by three factors: kernel latency, inter-device communication, and coordination overhead. PiKV addresses each of these challenges through specialized optimizations that work together to minimize end-to-end latency.
Figure 13 (top-left) shows the kernel latency breakdown for a 64K sequence. Routing dominates at long sequences (27-31% of total latency), a result of the need to evaluate gating logits for expert selection. Compression contributes 18-22% of latency, while scheduling and communication each account for 12-16% and 8-11% respectively.
The latency scaling with sequence length as shown in Figure 13 (top-right) demonstrates sublinear growth due to PiKV’s efficient caching and prefetching mechanisms. The latency increase from 1K to 64K sequences is only 3.2-3.8×, compared to 4.5-5.2× for traditional approaches.
Figure 13 (bottom-left) compares end-to-end latency across different configurations. PiKV achieves consistent performance across varying workload characteristics, while traditional approaches show significant variance (15-25%). The latency-performance trade-offs shown in Figure 13 (bottom-right) demonstrate that PiKV can achieve optimal performance across a wide range of accuracy requirements.
5.5. Throughput: Scaling with System Resources
Figure 14. Throughput analysis. Top-left: Area chart showing throughput scaling with GPU count with confidence bands and efficiency metrics; Top-right: Scatter plot of throughput-quality trade-offs with color-coded configurations and Pareto frontier; Bottom-left: Funnel chart showing throughput contribution by component with cumulative impact; Bottom-right: Candlestick chart of throughput over time showing system stability, variance, and trend analysis.
System throughput in distributed MoE inference is fundamentally limited by the coordination overhead between components and the mismatch between workload characteristics and resource availability. PiKV addresses this challenge through adaptive scaling and intelligent resource allocation.
Figure 14 (top-left) demonstrates PiKV’s near-linear scaling with GPU count, a result of its efficient resource allocation strategy that minimizes coordination overhead. Throughput scales with 94-97% efficiency as GPU count increases from 1 to 8, with only 3-6% overhead due to communication and synchronization.
The throughput scaling with different workload characteristics. Figure 14 (top-right) shows different configurations offering different throughput-quality trade-offs. High-throughput configurations like DistillationCompressor achieve 4.7× compression but with 12.2-13.0% accuracy degradation, while conservative configurations like LoRA++ achieve 2.8× compression with only 1.8-2.1% degradation.
Figure 14 (bottom-left) shows the throughput contribution by component. Routing contributes 23-27% of total throughput, compression contributes 33-37%, scheduling contributes 18-22%, and communication contributes 12-16%. The remaining 8-12% is attributed to system overhead and coordination.
5.6. Communication: Minimizing Distributed Overhead
Figure 15. Communication analysis. Top-left: Chord diagram showing communication patterns between GPU pairs with bandwidth indicators; Top-right: Network graph of communication overhead with edge weights indicating bandwidth and connection strength; Bottom-left: Gantt chart showing communication optimization impact over time with cumulative improvements; Bottom-right: Stream graph of communication bandwidth utilization showing temporal patterns and adaptive behavior.
Communication overhead is a critical bottleneck in distributed inference, particularly in MoE systems where expert routing patterns can create complex communication dependencies. PiKV addresses this challenge through intelligent communication patterns and optimization techniques.
Figure 15 (top-left) shows the distribution of communication patterns, with 58-62% being local patterns that require minimal inter-GPU communication. Remote patterns (28-32%) require single-hop communication, while broadcast patterns (8-12%) are used for synchronization and coordination.
Communication optimization is critical for maintaining high performance. Figure 15 (bottom-left) shows the impact of different optimization techniques, with predictive prefetching providing 23-27% improvement, compression providing 38-42% improvement, and async communication providing 15-19% improvement. The combined effect of all optimizations results in 45-52% overall improvement in communication efficiency.
Figure 15 (top-right) shows communication overhead between GPU pairs. The overhead varies significantly based on routing patterns and workload characteristics, with PiKV’s intelligent placement reducing average overhead by 32-38% compared to naive placement strategies. Figure 15 (bottom-right) demonstrates bandwidth utilization over time, showing that PiKV maintains high utilization (88-94%) while adapting to changing communication patterns.
5.7. Synthesis: Coordinated System Optimization
The comprehensive analysis reveals that PiKV achieves optimal performance through careful coordination of multiple system dimensions. The key insight is that no single optimization is sufficient—PiKV’s effectiveness stems from the synergistic interaction between routing, compression, and scheduling components.
The optimal configuration—AdaptiveRouter + PyramidCompressor + AdaKVScheduler—achieves 1.9× memory reduction, 82ms latency (20% improvement), 89% cache hit rate, and 1.0-1.2% accuracy degradation. This configuration represents a sweet spot that balances multiple competing objectives while maintaining system stability and reliability.
Performance scalability is equally important. PiKV demonstrates sublinear latency growth with sequence length, near-linear scaling with GPU count, and logarithmic scaling with expert count. These scaling characteristics ensure that PiKV can handle the growing demands of modern language models while maintaining predictable performance characteristics.
The unified design achieves 3.9× memory reduction over dense approaches, 1.7× latency improvement, and 2.7× throughput increase while maintaining minimal accuracy degradation. These results demonstrate that PiKV successfully addresses the system-level challenges of distributed MoE inference through coordinated optimization across multiple dimensions.
6. Conclusion
We present PiKV, a parallel and distributed KV cache management framework optimized for sparsely activated MoE-based large language models. PiKV introduces a KV cache management system for MoE, including sparse expert routing, cache compression, and stream-aware scheduling. This architecture rethinks KV caching not only as passive memory storage, but as a dynamic, query-driven retrieval system.
PiKV is a living project for scalable MoE serving, aiming for briding MoE sparsity and efficient system design optimization. Future work will explore online adaptation, hierarchical memory tiers, and integration with training-time sparsity strategies for end-to-end efficient large model deployment with MoE architecture.